### v1

# Dissipative currents and transport coefficients in relativistic spin hydrodynamics

Abstract
We determine the form of dissipative currents at the first order in relativistic spin hydrodynamics with finite chemical potential including gradients of the spin potential. Taking advantage of isotropy in the hydrodynamic local rest frame, using a suitable matching condition for the flow velocity and enforcing the semi-positivity of entropy production, we find 23 dissipative transport coefficients relating dissipative currents to gradients of the thermo-hydrodynamic fields: 4 for the symmetric part of the energy-momentum tensor, 5 for the antisymmetric part, 3 for the conserved vector current, and 11 for the spin tensor. We compare our finding with previous results in literature.
I Introduction
Relativistic spin hydrodynamics has emerged as a possible theory of phenomenological interest for the description of the strongly interacting Quark Gluon Plasma in relativistic heavy ion collisions, driven by experimental evidence of spin polarization of the produced particles  [1, 2]; for a recent review see, for example, Ref. [3]. For this reason, there has been a lot of interest in literature on this topic  [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], and yet its physical meaning is still under debate. In this theory, pseudo-gauge invariance of a combined transformation of energy-momentum and spin tensor is broken [38], and, unlike in the most usual formulation of relativistic hydrodynamics, the description of the fluid dynamics may need to include a non-vanishing spin tensor if, for instance, the system features a separation of scales in the relaxation times of spin and momentum degrees of freedom [10]. However, the relation between the spin tensor and the quantum field operators is not unique; in other words, the pseudo-gauge fixing remains arbitrary.
For any possible choice of the spin tensor operator, that is of the pseudo-gauge, the conservation equations read:
$\partial_{\mu}T^{\mu\nu}=0~{},~{}~{}\partial_{\mu}j^{\mu}=0~{},~{}~{}\partial_{\lambda}J^{\lambda\mu\nu}=0,$ (1)
where $T^{\mu\nu}$ is the mean value of the energy-momentum tensor, $j^{\mu}$ is the mean value of a conserved current, and $J^{\lambda\mu\nu}$ is the mean value of the total angular momentum tensor. The latter is decomposed into a so-called orbital part $L^{\lambda\mu\nu}=2T^{\lambda[\nu}x^{\mu]}$, [^1] We use square (round) brackets for antisymmetrization (symmetrization) of two operators, i.e., $A^{[\mu}B^{\nu]}=(A^{\mu}B^{\nu}-A^{\nu}B^{\mu})/2$ and $A^{(\mu}B^{\nu)}=(A^{\mu}B^{\nu}+A^{\nu}B^{\mu})/2$. and a spin tensor:
$J^{\lambda\mu\nu}=L^{\lambda\mu\nu}+S^{\lambda\mu\nu}.$ (2)
The conservation of total angular momentum implies:
$\partial_{\lambda}S^{\lambda\mu\nu}=T^{\nu\mu}-T^{\mu\nu}.$ (3)
A key problem in relativistic spin hydrodynamics is the derivation of dissipative currents as functions of gradients of thermo-hydrodynamic fields (four-velocity, temperature, chemical potentials, spin potential or suitable combinations thereof). Ambiguities may arise between different formulations depending on how these dissipative currents are decomposed, which can significantly impact the final physical results. In this work, we take advantage of a recent study [39] to provide general forms of dissipative currents up to first order in gradients of thermo-hydrodynamic fields. Our method is quite general and provides the maximal number of transport coefficients subject to the so-called matching condition of vanishing projections of the dissipative part of the currents onto the four-velocity field and the condition of positive entropy production.
The method can be summarized as follows: (i) in general, the dissipative currents are expressed as linear combinations of the gradients of thermo-hydrodynamic fields which appear in the entropy production rate expression, with tensor-valued coefficients; (ii) we impose that all tensor-valued coefficients are calculated at homogeneous global thermodynamic equilibrium such that they are invariant under spatial rotations in the local co-moving frame of the fluid. To ensure that this requirement is met, we decompose the tensor coefficients in terms of irreducible representations of the rotation group $\mathrm{SO}(3,\mathbb{R})$ and retain only the scalar part. (iii) finally, most of the remaining transport coefficients are eliminated by imposing matching conditions and the semi-positivity of the entropy production rate. These physical constraints reduce the 98 transport coefficients to 23, as tabulated in Sec. IV, thereby yielding the explicit forms of the various dissipative currents.
The paper is organized as follows. In Sec. II we show general expressions for dissipative currents and how to decompose them with the help of a set of irreducible bases. Then in Sec. III we impose the matching conditions and the semi-positivity constraints, which significantly reduces the number of unknown coefficients. The final results for dissipative currents and constraints for the corresponding transport coefficients are given in Sec. IV. In Sec. V, we verify a subset of these constraints for the specific microscopic framework of quantum kinetic theory. Finally, in Sec. VI we conclude our results with an outlook.
II Dissipative currents and their decomposition
In the quantum-statistical description of a relativistic fluid, the entropy production is obtained by calculating the divergence of the out-of-equilibrium entropy current, which is constructed from the local equilibrium density operator $\widehat{\rho}_{\rm LE}$, and the von Neumann entropy $S=-\mathrm{Tr}\big{[}\widehat{\rho}_{\rm LE}\log\widehat{\rho}_{\rm LE}\big{]}$. The resulting expression for the entropy production rate reads [39],
$\partial_{\mu}s^{\mu}=\left(T_{S}^{\mu\nu}-T^{\mu\nu}_{S,\,\rm LE}\right)\xi_{\mu\nu}-\left(j^{\mu}-j^{\mu}_{\rm LE}\right)\partial_{\mu}\zeta+\left(T^{\mu\nu}_{A}-T^{\mu\nu}_{A,\,\rm LE}\right)(\Omega_{\mu\nu}-\varpi_{\mu\nu})-\frac{1}{2}\left({\cal S}^{\mu\lambda\nu}-{\cal S}^{\mu\lambda\nu}_{\rm LE}\right)\partial_{\mu}\Omega_{\lambda\nu}\geq 0\,.$ (4)
Here, $\xi_{\mu\nu}=\partial_{(\mu}\beta_{\nu)}$ is the thermal shear tensor, $\varpi_{\mu\nu}=\partial_{[\nu}\beta_{\mu]}$ is the thermal vorticity tensor, and the fields $\beta_{\mu},~{}\zeta,~{}\Omega_{\mu\nu}$ are the thermal velocity four-vector $u_{\mu}/T$, the chemical potential over temperature $\mu/T$, and the spin potential over temperature $\omega_{\mu\nu}/T$, respectively.
In Eq. (4), quantities with subscript “LE” are mean values of the quantum operators at local equilibrium, which depend on the definition of a foliation, that is family of 3D hypersurfaces (see Fig. 1). Indeed, local equilibrium values of the currents are constrained to provide the exact physical values of the currents projected along $n^{\mu}$, the normal unit vector to the hypersurface, that is:
$n_{\mu}(\delta T^{\mu\nu}_{S}+\delta T^{\mu\nu}_{A})=0,~{}~{}~{}n_{\mu}\delta j^{\mu}=0,~{}~{}~{}n_{\mu}\delta S^{\mu\lambda\nu}=0\;,$ (5)
with $\delta X=X-X_{\text{LE}}$ for any quantity $X$. Provided that $n$ is a vorticity-free vector field (so that orthogonal hypersurfaces may exist)
$\epsilon^{\mu\nu\rho\sigma}n_{\nu}\partial_{\rho}n_{\sigma}=0\;,$ (6)
the choice of $n^{\mu}$ is arbitrary and corresponds to the freedom of choosing a frame in the classical formulation of relativistic hydrodynamics. A possible choice for $n^{\mu}$ is the fluid 4-velocity $u^{\mu}$ (as shown in Fig. 1) motivated by the fact that with this choice one recovers matching conditions similar to those in Landau’s or Eckart’s formulations of dissipative hydrodynamics without a spin tensor. However, in the quantum relativistic statistical formulation which led to Eq. (4), the condition (6) applied to the four-velocity field implies a restriction which is absent in the classical formulation and should be kept in mind when comparing the final results. Particularly, if the velocity field has a finite vorticity (like at global thermodynamic equilibrium with non-vanishing angular velocity), $n$ cannot coincide with $u$ and should be defined otherwise (see the discussion in Ref. [40]).
Figure 1: Three-dimensional foliation of initial Cauchy space-like hyperspace $\Sigma_{(0)}$, with volume $\Omega$, and present time hypersurface $\Sigma_{(t)}$, where $\hat{n}$ is chosen to be the fluid four-velocity.
Equation (4) makes it clear which thermo-hydrodynamic gradients contribute to entropy production. Therefore, we identify these contributions - in the usual classification scheme - as being of first order in gradients:
$\xi_{\mu\nu}\sim\mathcal{O}(\partial)~{},~{}~{}\partial_{\mu}\zeta\sim\mathcal{O}(\partial)~{},~{}~{}(\Omega_{\mu\nu}-\varpi_{\mu\nu})\sim\mathcal{O}(\partial)~{},~{}~{}\partial_{\mu}\Omega_{\lambda\nu}\sim\mathcal{O}(\partial),$ (7)
and we will use them to express the dissipative currents at first order in the gradient expansion. It should be noted that in other studies the spin potential $\Omega$ (and the thermal vorticity as well) is itself considered as ”first-order quantity” in the gradients so that $\partial\Omega$ is neglected as being a ”second order quantity”. We do not make such an assumption, thus we tacitly assume $\Omega$ and $\varpi$ to be not necessarily small. On the other hand, we will neglect the anisotropy induced by a potentially large thermal vorticity and spin potential when decomposing the tensor-valued coefficients, as will be made clear shortly.
At first order in gradients we can write in general:
$\delta T^{\mu\nu}_{S}=H^{\mu\nu\rho\sigma}\xi_{\rho\sigma}+\frac{1}{T}K^{\mu\nu\rho}\partial_{\rho}\zeta+L^{\mu\nu\rho\sigma}\left(\Omega_{\rho\sigma}-\varpi_{\rho\sigma}\right)+\frac{1}{T}M^{\mu\nu\rho\sigma\tau}\partial_{\rho}\Omega_{\sigma\tau}\,,$ (8)
$\delta T^{\mu\nu}_{A}=N^{\mu\nu\rho\sigma}\xi_{\rho\sigma}+\frac{1}{T}P^{\mu\nu\rho}\partial_{\rho}\zeta+Q^{\mu\nu\rho\sigma}\left(\Omega_{\rho\sigma}-\varpi_{\rho\sigma}\right)+\frac{1}{T}R^{\mu\nu\rho\sigma\tau}\partial_{\rho}\Omega_{\sigma\tau}\,,$ (9)
$T\delta j^{\mu}=G^{\mu\rho\sigma}\xi_{\rho\sigma}+\frac{1}{T}I^{\mu\rho}\partial_{\rho}\zeta+O^{\mu\rho\sigma}(\Omega_{\rho\sigma}-\varpi_{\rho\sigma})+\frac{1}{T}F^{\mu\rho\sigma\tau}\partial_{\rho}\Omega_{\sigma\tau}\,,$ (10)
$T\delta S^{\mu\lambda\nu}=T^{\mu\lambda\nu\rho\sigma}\xi_{\rho\sigma}+\frac{1}{T}U^{\mu\lambda\nu\rho}\partial_{\rho}\zeta+V^{\mu\lambda\nu\rho\sigma}\left(\Omega_{\rho\sigma}-\varpi_{\rho\sigma}\right)+\frac{1}{T}W^{\mu\lambda\nu\rho\sigma\tau}\partial_{\rho}\Omega_{\sigma\tau}\,,$ (11)
where we have introduced tensor coefficients $H,K,L,M,\dots,W$ with suitable rank and symmetry in the exchange of indices; the additional factors $T$ and $1/T$ are introduced such that all the tensor coefficients have the same dimension.
The tensor coefficients are generally dependent on the quantities that characterize global equilibrium. In general, in Minkowski spacetime, this depends on 10 parameters (plus a set of possible chemical potentials): a constant time-like vector and an antisymmetric constant tensor, the thermal vorticity $\varpi$. The possible dependencies of the tensor coefficients in (8)-(11) on those parameters can be, in general, very complicated. In practice, the absence of symmetries of the density operator in the most general form of global thermodynamic equilibrium with $\varpi\neq 0$ makes the number of relevant scalar coefficients equal to the number of independent components of the tensors. However, if we neglect the dependence of those tensor coefficients on thermal vorticity $\varpi$ and we expand them from homogeneous global equilibrium, where the four-temperature $\beta$ is constant and $\varpi=0$, they must be invariant under spatial rotations in the local co-moving frame of the fluid, and this constraint greatly reduces the number of relevant scalar coefficients. To fulfill this requirement, we decompose the tensor coefficients in terms of irreducible representations of the rotation group ${\rm SO}(3,\mathbb{R})$. The irreducible basis is built on the fluid 4-velocity $u^{\mu}$, the transverse projector $\Delta^{\mu\nu}\equiv g^{\mu\nu}-u^{\mu}u^{\nu}$, and the rank-3 tensor $\epsilon^{\mu\nu\lambda\gamma}u_{\gamma}$, which are rotationally invariant in the rest frame associated to $u^{\mu}$. The details on the approach for constructing this basis are discussed in Appendix A.
In Tables 1-5, we list all the irreducible bases and the corresponding factors for each of the tensor coefficients $H^{\mu\nu\rho\sigma},K^{\mu\nu\rho},\dots,W^{\mu\lambda\nu\rho\sigma\tau}$, where the factors are denoted by lowercase Latin letters with numbers as lower indices. For example, from Table 1 we can read out the irreducible decomposition of $H^{\mu\nu\rho\sigma}$ as
$H^{\mu\nu\rho\sigma}=h_{1}u^{(\mu}\Delta^{\nu)(\rho}u^{\sigma)}+h_{2}\Delta^{\mu\nu}\Delta^{\rho\sigma}+h_{3}\Delta^{\mu\nu,\rho\sigma}+h_{4}\Delta^{\mu\nu}u^{\rho}u^{\sigma}+h_{5}u^{\mu}u^{\nu}\Delta^{\rho\sigma}+h_{6}u^{\mu}u^{\nu}u^{\rho}u^{\sigma}\;,$ (12)
where the traceless symmetric projector $\Delta_{\mu\nu,\alpha\beta}\equiv(\Delta_{\mu\alpha}\Delta_{\nu\beta}+\Delta_{\nu\alpha}\Delta_{\mu\beta})/2-\Delta_{\mu\nu}\Delta_{\alpha\beta}/3$ was used. Substituting the decomposed tensor coefficients $H,K,L,M,\dots,W$ back into the dissipative currents in Eqs. (8)-(11), and subsequently into the entropy production rate (4), results in an explicit expression of $\partial_{\mu}s^{\mu}$, which will be analyzed further in the following.
$H^{\mu\nu\rho\sigma}$ $K^{\mu\nu\rho}$ $L^{\mu\nu\rho\sigma}$ $M^{\mu\nu\rho\sigma\tau}$
$h_{1}$ $u^{(\mu}\Delta^{\nu)(\rho}u^{\sigma)}$ $k_{1}$ $u^{(\mu}\Delta^{\nu)\rho}$ $l_{1}$ $u^{(\mu}\Delta^{\nu)[\rho}u^{\sigma]}$ $m_{1}$ $u^{(\mu}\Delta^{\nu)[\sigma}u^{\tau]}u^{\rho}$ $m_{7}$ $u^{\mu}u^{\nu}u^{[\sigma}\Delta^{\tau]\rho}$
$h_{2}$ $\Delta^{\mu\nu}\Delta^{\rho\sigma}$ $k_{2}$ $\Delta^{\mu\nu}u^{\rho}$ $l_{2}$ $u^{(\mu}\epsilon^{\nu)\rho\sigma\gamma}u_{\gamma}$ $m_{2}$ $u^{(\mu}\epsilon^{\nu)\sigma\tau\gamma}u_{\gamma}u^{\rho}$ $m_{8}$ $u^{\mu}u^{\nu}\epsilon^{\sigma\tau\gamma\rho}u_{\gamma}$
$h_{3}$ $\Delta^{\mu\nu,\rho\sigma}$ $k_{3}$ $u^{\mu}u^{\nu}u^{\rho}$ $m_{3}$ $\Delta^{\mu\nu}u^{[\sigma}\Delta^{\tau]\rho}$ $m_{9}$ $(\epsilon^{\sigma\tau\gamma(\mu}\Delta^{\nu)\rho}+\frac{1}{3}\Delta^{\mu\nu}\epsilon^{\rho\sigma\tau\gamma})u_{\gamma}$
$h_{4}$ $\Delta^{\mu\nu}u^{\rho}u^{\sigma}$ $m_{4}$ $\Delta^{\mu\nu}\epsilon^{\sigma\tau\gamma\rho}u_{\gamma}$ $m_{10}$ $\Delta^{\mu\nu,\rho[\tau}u^{\sigma]}$
$h_{5}$ $u^{\mu}u^{\nu}\Delta^{\rho\sigma}$ $m_{5}$ $2u^{(\mu}\Delta^{\nu)[\sigma}\Delta^{\tau]\rho}$
$h_{6}$ $u^{\mu}u^{\nu}u^{\rho}u^{\sigma}$ $m_{6}$ $u^{(\mu}\epsilon^{\nu)\rho\gamma[\sigma}u^{\tau]}u_{\gamma}$
Table 1: Irreducible decompositions of $H^{\mu\nu\rho\sigma}$, $K^{\mu\nu\rho}$, $L^{\mu\nu\rho\sigma}$, and $M^{\mu\nu\rho\sigma\tau}$.
$N^{\mu\nu\rho\sigma}$ $P^{\mu\nu\rho}$ $Q^{\mu\nu\rho\sigma}$ $R^{\mu\nu\rho\sigma\tau}$
$n_{1}$ $u^{[\mu}\Delta^{\nu](\rho}u^{\sigma)}$ $p_{1}$ $\Delta^{\rho[\mu}u^{\nu]}$ $q_{1}$ $u^{[\mu}\Delta^{\nu][\rho}u^{\sigma]}$ $r_{1}$ $u^{[\mu}\Delta^{\nu][\sigma}u^{\tau]}u^{\rho}$ $r_{5}$ $\epsilon^{\mu\nu\gamma[\sigma}u^{\tau]}u_{\gamma}u^{\rho}$
[…]
$q_{4}$ $u^{[\mu}\epsilon^{\nu]\rho\sigma\gamma}u_{\gamma}$ $r_{4}$ $u^{[\mu}\epsilon^{\nu]\gamma\rho[\sigma}u^{\tau]}u_{\gamma}$ $r_{8}$ $2\Delta^{\rho[\mu}\Delta^{\nu][\sigma}u^{\tau]}$
Table 2: Irreducible decompositions of $N^{\mu\nu\rho\sigma}$, $P^{\mu\nu\rho}$, $Q^{\mu\nu\rho\sigma}$, and $R^{\mu\nu\rho\sigma\tau}$.
$G^{\mu\rho\sigma}$ $I^{\mu\rho}$ $O^{\mu\rho\sigma}$ $F^{\mu\rho\sigma\tau}$
$g_{1}$ $\Delta^{\mu(\rho}u^{\sigma)}$ $i_{1}$ $\Delta^{\mu\rho}$ $o_{1}$ $\Delta^{\mu[\rho}u^{\sigma]}$ $f_{1}$ $\Delta^{\mu[\sigma}u^{\tau]}u^{\rho}$ $f_{4}$ $\epsilon^{\mu\rho\gamma[\sigma}u^{\tau]}u_{\gamma}$
$g_{2}$ $u^{\mu}\Delta^{\rho\sigma}$ $i_{2}$ $u^{\mu}u^{\rho}$ $o_{2}$ $\epsilon^{\mu\rho\sigma\gamma}u_{\gamma}$ $f_{2}$ $\epsilon^{\mu\sigma\tau\gamma}u^{\rho}u_{\gamma}$ $f_{5}$ $u^{\mu}\Delta^{\rho[\sigma}u^{\tau]}$
$g_{3}$ $u^{\mu}u^{\rho}u^{\sigma}$ $f_{3}$ $2\Delta^{\mu[\sigma}\Delta^{\tau]\rho}$ $f_{6}$ $u^{\mu}\epsilon^{\rho\sigma\tau\gamma}u_{\gamma}$
Table 3: Irreducible decompositions of $G^{\mu\rho\sigma}$, $I^{\mu\rho}$, $O^{\mu\rho\sigma}$, and $F^{\mu\rho\sigma\tau}$.
$T^{\mu\lambda\nu\rho\sigma}$ $U^{\mu\lambda\nu\rho}$ $V^{\mu\lambda\nu\rho\sigma}$
$t_{1}$ $\Delta^{\mu[\nu}u^{\lambda]}u^{\rho}u^{\sigma}$ $t_{6}$ $u^{[\lambda}\epsilon^{\nu]\mu\gamma(\rho}u^{\sigma)}u_{\gamma}$ $u_{1}$ $\Delta^{\mu[\nu}u^{\lambda]}u^{\rho}$ $v_{1}$ $2\Delta^{\mu[\rho}\Delta^{\sigma][\nu}u^{\lambda]}$
$t_{2}$ $\Delta^{\mu[\nu}u^{\lambda]}\Delta^{\rho\sigma}$ $t_{7}$ $u^{[\lambda}\Delta^{\nu]\mu,\rho\sigma}$ $u_{2}$ $2\Delta^{\mu[\nu}\Delta^{\lambda]\rho}$ $v_{2}$ $2\Delta^{\mu[\rho}\epsilon^{\sigma]\nu\lambda\gamma}u_{\gamma}$
$t_{3}$ $2\Delta^{\mu[\nu}\Delta^{\lambda](\rho}u^{\sigma)}$ $t_{8}$ $(\Delta^{\mu(\rho}\epsilon^{\sigma)\nu\lambda\gamma}-\Delta^{\rho\sigma}\epsilon^{\mu\nu\lambda\gamma}/3)u_{\gamma}$ $u_{3}$ $\epsilon^{\mu\nu\lambda\gamma}u_{\gamma}u^{\rho}$ $v_{3}$ $2\Delta^{\mu[\lambda}\Delta^{\nu][\rho}u^{\sigma]}$
$t_{4}$ $\epsilon^{\mu\nu\lambda\gamma}u_{\gamma}u^{\rho}u^{\sigma}$ $t_{9}$ $u^{\mu}u^{[\lambda}\Delta^{\nu](\rho}u^{\sigma)}$ $u_{4}$ $\epsilon^{\mu\rho\gamma[\nu}u^{\lambda]}u_{\gamma}$ $v_{4}$ $u^{[\rho}\epsilon^{\sigma]\mu\gamma[\lambda}u^{\nu]}u_{\gamma}$
$t_{5}$ $\epsilon^{\mu\nu\lambda\gamma}u_{\gamma}\Delta^{\rho\sigma}$ $t_{10}$ $u^{\mu}\epsilon^{\lambda\nu\gamma(\rho}u^{\sigma)}u_{\gamma}$ $u_{5}$ $u^{\mu}\Delta^{\rho[\nu}u^{\lambda]}$ $v_{5}$ $u^{\mu}u^{[\lambda}\Delta^{\nu][\rho}u^{\sigma]}$
$u_{6}$ $u^{\mu}\epsilon^{\nu\lambda\rho\gamma}u_{\gamma}$ $v_{6}$ $u^{\mu}u^{[\nu}\epsilon^{\lambda]\rho\sigma\gamma}u_{\gamma}$
$v_{7}$ $u^{\mu}\epsilon^{\nu\lambda\gamma[\rho}u^{\sigma]}u_{\gamma}$
$v_{8}$ $u^{\mu}\Delta^{\lambda[\rho}\Delta^{\sigma]\nu}$
Table 4: Irreducible decompositions of $T^{\mu\lambda\nu\rho\sigma}$, $U^{\mu\lambda\nu\rho}$, and $V^{\mu\lambda\nu\rho\sigma}$.
$W^{\mu\lambda\nu\rho\sigma\tau}$
$w_{1}$ $\Delta^{\mu[\nu}u^{\lambda]}u^{[\sigma}\Delta^{\tau]\rho}$ $w_{7}$ $2\Delta^{\mu[\nu}\Delta^{\lambda][\tau}u^{\sigma]}u^{\rho}$ $w_{13}$ $u^{\mu}u^{\rho}\Delta^{\lambda[\sigma}\Delta^{\tau]\nu}$
$w_{2}$ $\epsilon^{\mu\lambda\nu\gamma}u_{\gamma}\epsilon^{\rho\sigma\tau\kappa}u_{\kappa}$ $w_{8}$ $2\Delta^{\mu[\tau}\epsilon^{\sigma]\lambda\nu\gamma}u_{\gamma}u^{\rho}$ $w_{14}$ $u^{\mu}\epsilon^{\lambda\nu\gamma[\sigma}u^{\tau]}u^{\rho}u_{\gamma}$
$w_{3}$ $\Delta^{\mu[\nu}u^{\lambda]}\epsilon^{\sigma\tau\gamma\rho}u_{\gamma}+\epsilon^{\mu\lambda\nu\gamma}u_{\gamma}\Delta^{\rho[\sigma}u^{\tau]}$ $w_{9}$ $4\Delta^{\mu[\lambda}\Delta^{\nu][\sigma}\Delta^{\tau]\rho}$ $w_{15}$ $u^{\mu}u^{[\nu}\epsilon^{\lambda]\rho\gamma[\sigma}u^{\tau]}u_{\gamma}$
$w_{4}$ $\Delta^{\mu[\nu}u^{\lambda]}\epsilon^{\sigma\tau\gamma\rho}u_{\gamma}-\epsilon^{\mu\lambda\nu\gamma}u_{\gamma}\Delta^{\rho[\sigma}u^{\tau]}$ $w_{10}$ $u^{[\lambda}\Delta^{\nu]\mu,\rho\tau}u^{\sigma]}$ $w_{16}$ $u^{\mu}\Delta^{\rho[\nu}\Delta^{\lambda][\sigma}u^{\tau]}$
$w_{5}$ $u^{[\nu}\epsilon^{\lambda]\mu\gamma[\sigma}u^{\tau]}u_{\gamma}u^{\rho}$ $w_{11}$ $u^{\mu}u^{[\lambda}\Delta^{\nu][\sigma}u^{\tau]}u^{\rho}$ $w_{17}$ $u^{\mu}u^{[\nu}\Delta^{\lambda][\tau}\Delta^{\sigma]\rho}$
$w_{6}$ $2\Delta^{\mu[\sigma}\Delta^{\tau][\nu}u^{\lambda]}u^{\rho}$ $w_{12}$ $u^{\mu}u^{[\lambda}\epsilon^{\nu]\sigma\tau\gamma}u_{\gamma}u^{\rho}$ $w_{18}$ $u^{\mu}\Delta^{\rho[\sigma}\epsilon^{\tau]\lambda\nu\gamma}u_{\gamma}$
$w_{19}$ $\Delta^{\mu\rho}u^{[\tau}\Delta^{\sigma][\lambda}u^{\nu]}-\Delta^{\mu[\tau}u^{\sigma]}u^{[\lambda}\Delta^{\nu]\rho}$
$w_{20}$ $\Delta^{\mu[\nu}\epsilon^{\lambda]\rho\gamma[\tau}u^{\sigma]}u_{\gamma}+(\Delta^{\rho[\lambda}u^{\nu]}\epsilon^{\mu\sigma\tau\gamma}+\Delta^{\mu\rho}u^{[\lambda}\epsilon^{\nu]\sigma\tau\gamma})u_{\gamma}/2$
$w_{21}$ $\Delta^{\mu[\nu}\epsilon^{\lambda]\rho\gamma[\tau}u^{\sigma]}u_{\gamma}-(\Delta^{\rho[\lambda}u^{\nu]}\epsilon^{\mu\sigma\tau\gamma}+\Delta^{\mu\rho}u^{[\lambda}\epsilon^{\nu]\sigma\tau\gamma})u_{\gamma}/2$
$w_{22}$ $\Delta^{\mu\rho}\epsilon^{\nu\lambda\gamma[\sigma}u^{\tau]}u_{\gamma}-\Delta^{\mu[\nu}\epsilon^{\lambda]\rho\gamma[\tau}u^{\sigma]}u_{\gamma}-\epsilon^{\mu\nu\lambda\gamma}u_{\gamma}u^{[\sigma}\Delta^{\tau]\rho}/3$
$w_{23}$ $(\Delta^{\rho[\lambda}u^{\nu]}\epsilon^{\mu\sigma\tau\gamma}-\Delta^{\mu\rho}u^{[\lambda}\epsilon^{\nu]\sigma\tau\gamma})u_{\gamma}/2-\Delta^{\mu[\nu}u^{\lambda]}\epsilon^{\sigma\tau\gamma\rho}u_{\gamma}/3$
$w_{24}$ $\Delta^{\mu\rho}\Delta^{\lambda[\sigma}\Delta^{\tau]\nu}+\epsilon^{\mu\lambda\nu\gamma}u_{\gamma}\epsilon^{\rho\sigma\tau\kappa}u_{\kappa}/6-\Delta^{\mu[\nu}\Delta^{\lambda][\sigma}\Delta^{\tau]\rho}$
Table 5: Irreducible decompositions of $W^{\mu\lambda\nu\rho\sigma\tau}$.
III Matching conditions and semi-positivity constraints
Tables 15 introduce 98 undetermined factors, leading to very complicated forms of the dissipative currents. Fortunately, some of these factors can be eliminated by imposing matching conditions and the semi-positivity of the entropy production rate $\partial_{\mu}s^{\mu}\geq 0$. In this work, we impose the following matching conditions:
$u_{\mu}(\delta T^{\mu\nu}_{S}+\delta T^{\mu\nu}_{A})=0,~{}~{}~{}u_{\mu}\delta j^{\mu}=0,~{}~{}~{}u_{\mu}\delta S^{\mu\lambda\nu}=0.$ (13)
By substituting Eqs. (8)-(11) into (13) and applying the irreducible decompositions in Tables 1-5, the conditions are converted to,
$u_{\mu}\left[\delta T^{\mu\nu}_{S}+\delta T^{\mu\nu}_{A}\right]=0$$\quad\implies\quad\begin{cases}&h_{1}+n_{1}=0,\quad h_{5}=h_{6}=0,\quad k_{1}-p_{1}=0,\quad k_{3}=0,\\ &l_{1}+q_{1}=0,\quad l_{2}+{q_{4}}=0,\quad m_{1}+r_{1}=0,\quad m_{2}+r_{2}=0,\\ &m_{5}+r_{3}=0,\quad m_{6}-r_{4}=0,\quad m_{7}=m_{8}=0.\end{cases}$ (14)
$u_{\mu}\delta j^{\mu}=0$ $\quad\implies\quad\hskip 18.49411ptg_{2}=g_{3}=0,\quad i_{2}=0,\quad f_{5}=f_{6}=0.$ (15)
$u_{\mu}\delta S^{\mu\lambda\nu}=0$$\quad\implies\quad\begin{cases}&t_{9}=t_{10}=0,\quad u_{5}=u_{6}=0,\quad v_{5}=v_{6}=0,\quad v_{7}=v_{8}=0,\\ &w_{11}=w_{12}=0,\quad w_{13}=w_{14}=0,\quad w_{15}=w_{16}=0,\quad w_{17}=w_{18}=0.\end{cases}$ (16)
After taking these 34 constraints into account, we are left with $98-34=64$ undetermined coefficients.
On the other hand, the second law of thermodynamics requires the semi-positivity of entropy production rate, given in Eq. (4). In order to simplify this condition, we decompose the thermo-hydrodynamic gradient terms using the irreducible representations of the rotation group, identical to the procedure discussed in the preceding section that led to Tables 1-5. We start from the symmetric shear tensor,
$\xi^{\mu\nu}=(D\beta)u^{\mu}u^{\nu}+\frac{1}{3T}\theta\Delta^{\mu\nu}+\frac{1}{T}\mathcal{J}_{h}^{\alpha}\left(\Delta^{\mu}_{\alpha}u^{\nu}+\Delta^{\nu}_{\alpha}u^{\mu}\right)+\frac{1}{T}\sigma^{\alpha\beta}\Delta^{\mu\nu}_{\alpha\beta},$ (17)
where we have introduced the comoving and spatial derivatives $D\equiv u_{\mu}\partial^{\mu}$ and $\nabla^{\mu}\equiv\Delta^{\mu\nu}\partial_{\nu}$, the expansion scalar $\theta\equiv\partial_{\mu}u^{\mu}$, the heat flow $\mathcal{J}_{h}^{\mu}\equiv Du^{\mu}-(1/T)\nabla^{\mu}T$, and the shear tensor $\sigma^{\mu\nu}\equiv\Delta^{\mu\nu}_{\alpha\beta}\partial^{\alpha}u^{\beta}$. The following gradient term, namely the difference between the spin potential and the thermal vorticity, $\Omega_{\mu\nu}-\varpi_{\mu\nu}$, is antisymmetric and can be decomposed as,
$\Omega^{\mu\nu}-\varpi^{\mu\nu}=\frac{1}{T}\mathcal{E}^{\alpha}\left(\Delta^{\nu}_{\alpha}u^{\mu}-\Delta^{\mu}_{\alpha}u^{\nu}\right)-\frac{1}{2T}\epsilon^{\mu\nu\alpha\beta}u_{\alpha}\mathcal{B}_{\beta},$ (18)
where the electric-like part is $\mathcal{E}^{\mu}\equiv Tu_{\nu}(\Omega^{\nu\mu}-\varpi^{\nu\mu})$, and the magnetic-like part reads $\mathcal{B}^{\mu}\equiv T\epsilon^{\alpha\beta\gamma\mu}u_{\gamma}(\Omega_{\alpha\beta}-\varpi_{\alpha\beta})$. The decomposition of the next gradient contribution, i.e., the derivative of the of the chemical potential over temperature, is given by
$\partial_{\mu}\zeta=\left(D\zeta\right)u_{\mu}+\left(\nabla_{\alpha}\zeta\right)\Delta^{\alpha}_{\mu}\;.$ (19)
Finally, the derivative of the spin potential is decomposed as follows, with details of the derivation relegated to Appendix A,
$\partial_{\mu}\Omega_{\lambda\nu}=$ $\ 2\mathcal{X}^{\gamma}u_{\mu}u_{[\lambda}\Delta_{\nu]\gamma}-\frac{1}{2}\mathcal{Y}^{\gamma}u_{\mu}\epsilon_{\lambda\nu\sigma\gamma}u^{\sigma}+\frac{1}{2}\mathcal{Z}u_{[\lambda}\Delta_{\nu]\mu}+\mathcal{T}^{\gamma}u_{[\lambda}\epsilon_{\nu]\mu\alpha\gamma}u^{\alpha}+2\mathcal{F}^{\rho\sigma}u_{[\lambda}\Delta_{\nu]\sigma}\Delta_{\mu\rho}$
$\ -\frac{1}{8}\mathcal{H}\epsilon_{\lambda\nu\alpha\mu}u^{\alpha}-\frac{1}{2}\mathcal{G}_{[\lambda}\Delta_{\nu]\mu}-\frac{1}{2}\mathcal{I}^{\rho\sigma}\Delta_{\mu\rho}\epsilon_{\lambda\nu\tau\sigma}u^{\tau}\;.$ (20)
The scalars, vectors, and tensors appearing in the expression above are defined as follows,
$\mathcal{X}^{\gamma}=u^{[\rho}\Delta^{\sigma]\gamma}D\Omega_{\rho\sigma},\quad\mathcal{Y}^{\gamma}=\epsilon^{\rho\sigma\tau\gamma}u_{\tau}D\Omega_{\rho\sigma},\quad\mathcal{Z}=u^{[\rho}\nabla^{\sigma]}\Omega_{\rho\sigma},\quad\mathcal{T}^{\gamma}=u^{[\rho}\epsilon^{\sigma]\lambda\theta\gamma}u_{\lambda}\partial_{\theta}\Omega_{\rho\sigma},$
$\mathcal{F}^{\rho\sigma}=u^{[\gamma}\Delta^{\theta](\rho}\nabla^{\sigma)}\Omega_{\gamma\theta}-\Delta^{\rho\sigma}\mathcal{Z}/4,\quad\mathcal{H}=\epsilon^{\rho\sigma\lambda\gamma}u_{\lambda}\partial_{\gamma}\Omega_{\rho\sigma},\quad\mathcal{G}^{\gamma}=2\Delta^{\gamma\tau}\nabla^{\rho}\Omega_{\rho\tau},$
$\mathcal{I}^{\rho\sigma}=u_{\gamma}\epsilon^{\theta\lambda\gamma(\rho}\nabla^{\sigma)}\Omega_{\theta\lambda}-\Delta^{\rho\sigma}\mathcal{H}/4.$ (21)
Note that the tensors $\mathcal{F}^{\mu\nu}$ and $\mathcal{I}^{\mu\nu}$ are symmetric, but not traceless, with their traces being related to $\mathcal{Z}$ and $\mathcal{H}$, respectively. They can be further written in terms of traceless symmetric tensors $\mathcal{F}^{\mu\nu}_{S}$ and $\mathcal{I}^{\mu\nu}_{S}$ as
$\mathcal{F}^{\mu\nu}=\mathcal{F}^{\mu\nu}_{S}+\frac{1}{12}\Delta^{\mu\nu}\mathcal{Z},\quad\mathcal{I}^{\mu\nu}=\mathcal{I}^{\mu\nu}_{S}+\frac{1}{12}\Delta^{\mu\nu}\mathcal{H}.$ (22)
Then by contracting the above gradient terms (17)-(III) with the dissipative currents (8)–(11), and applying the matching conditions (14)-(16), we arrive at the following results
$T^{2}\xi_{\mu\nu}\delta T^{\mu\nu}_{S}=\,h_{1}\mathcal{J}_{h}^{\mu}\mathcal{J}_{h,\mu}+h_{2}\theta^{2}+h_{3}\sigma^{\mu\nu}\sigma_{\mu\nu}+\left(h_{4}TD\beta+k_{2}D\zeta+m_{3}\mathcal{Z}+m_{4}\mathcal{H}\right)\theta$
$\hskip 59.75095pt-\mathcal{J}^{h}_{\mu}\left(-k_{1}\nabla^{\mu}\zeta+l_{1}\mathcal{E}^{\mu}+l_{2}\mathcal{B}^{\mu}+m_{1}\mathcal{X}^{\mu}+m_{2}\mathcal{Y}^{\mu}+m_{5}\mathcal{G}^{\mu}+m_{6}\mathcal{T}^{\mu}\right)$
$\hskip 59.75095pt+\sigma_{\mu\nu}\left(m_{9}\mathcal{I}_{S}^{\mu\nu}+m_{10}\mathcal{F}_{S}^{\mu\nu}\right)\,,$ (23)
$T^{2}\partial_{\mu}\zeta\delta j^{\mu}=\,i_{1}(\nabla^{\mu}\zeta)(\nabla_{\mu}\zeta)-\left(-g_{1}\mathcal{J}_{h}^{\mu}+o_{1}\mathcal{E}^{\mu}+o_{2}\mathcal{B}^{\mu}+f_{1}\mathcal{X}^{\mu}+f_{2}\mathcal{Y}^{\mu}+f_{3}\mathcal{G}^{\mu}+f_{4}\mathcal{T}^{\mu}\right)\nabla_{\mu}\zeta\,,$ (24)
$T^{2}(\Omega_{\mu\nu}-\varpi_{\mu\nu})\delta T^{\mu\nu}_{A}=\,-q_{1}\mathcal{E}^{\mu}\mathcal{E}_{\mu}-q_{2}\mathcal{B}^{\mu}\mathcal{B}_{\mu}-(q_{3}+q_{4})\mathcal{E}^{\mu}\mathcal{B}_{\mu}$
$\hskip 102.43008pt-\mathcal{E}_{\mu}\left(-n_{1}\mathcal{J}_{h}^{\mu}+p_{1}\nabla^{\mu}\zeta+r_{1}\mathcal{X}^{\mu}+r_{2}\mathcal{Y}^{\mu}+r_{3}\mathcal{G}^{\mu}-r_{4}\mathcal{T}^{\mu}\right)$
$\hskip 102.43008pt-\mathcal{B}_{\mu}\left(-n_{2}\mathcal{J}_{h}^{\mu}+p_{2}\nabla^{\mu}\zeta+r_{5}\mathcal{X}^{\mu}+r_{6}\mathcal{Y}^{\mu}+r_{7}\mathcal{G}^{\mu}-r_{8}\mathcal{T}^{\mu}\right)\,,$ (25)
$T^{2}\partial_{\mu}\Omega_{\lambda\nu}\delta S^{\mu\lambda\nu}=\,(t_{1}TD\beta+t_{2}\theta+u_{1}D\zeta)\mathcal{Z}+(t_{4}TD\beta+t_{5}\theta+u_{3}D\zeta)\mathcal{H}+w_{1}\mathcal{Z}^{2}+w_{2}\mathcal{H}^{2}+2w_{3}\mathcal{Z}\mathcal{H}$
$\hskip 76.82234pt-(t_{3}\mathcal{J}^{h}_{\mu}+u_{2}\nabla_{\mu}\zeta+v_{2}\mathcal{B}_{\mu}+v_{3}\mathcal{E}_{\mu}+w_{7}\mathcal{X_{\mu}}+w_{8}\mathcal{Y}_{\mu})\mathcal{G}^{\mu}$
$\hskip 76.82234pt-(t_{6}\mathcal{J}^{h}_{\mu}+u_{4}\nabla_{\mu}\zeta+v_{1}\mathcal{B}_{\mu}+v_{4}\mathcal{E}_{\mu}+w_{5}\mathcal{X_{\mu}}+w_{6}\mathcal{Y}_{\mu})\mathcal{T}^{\mu}$
$\hskip 76.82234pt-w_{9}\mathcal{G}^{\mu}\mathcal{G}_{\mu}-w_{19}\mathcal{T}^{\mu}\mathcal{T}_{\mu}-w_{21}\mathcal{G}^{\mu}\mathcal{T}_{\mu}$
$\hskip 76.82234pt+(t_{7}\mathcal{F}^{\mu\nu}_{S}+t_{8}I^{\mu\nu}_{S})\sigma_{\mu\nu}+w_{10}\mathcal{F}^{\mu\nu}_{S}\mathcal{F}_{S,\mu\nu}+(w_{22}+w_{23})\mathcal{I}^{\mu\nu}_{S}\mathcal{F}_{S,\mu\nu}-\frac{1}{2}w_{24}\mathcal{I}^{\mu\nu}_{S}\mathcal{I}_{S,\mu\nu}\,.$ (26)
The total entropy production rate, $\partial_{\mu}s^{\mu}$, is then obtained by substituting these results into Eq. (4).
Scalars Pseudo-scalars Vectors Pseudo-vectors Tensors Pseudo-tensors
[…]
$TD\beta$ —- $\nabla^{\mu}\zeta$ $\mathcal{Y}^{\mu}$ $\mathcal{F}^{\mu\nu}_{S}$
$D\zeta$ —- $\mathcal{X}^{\mu}$ $\mathcal{T}^{\mu}$ —-
$\mathcal{Z}$ —- $\mathcal{E}^{\mu}$
—– $\mathcal{G}^{\mu}$
Table 6: Classification of the thermo-hydrodynamic gradient terms based on their properties under reflections and rotations.
We note that $\partial_{\mu}s^{\mu}$ is a Lorentz scalar which is invariant under reflections as well as spatial rotations in the rest frame associated to $u^{\mu}$. Regarding the thermo-hydrodynamic gradient terms appearing in Eqs. (III)-(26), we may classify them according to their behavior under this class of transformations, as listed in Table 6. In order to ensure all of the terms appearing in the entropy production rate are (parity-even) scalars, some of the coefficients in Eqs. (III)-(26) have to be parity-odd while others are parity-even. For simplicity, we set all parity-odd coefficients to zero. This further eliminates 28 of the coefficients, bringing down the total number to $64-28=36$. Then, we can use the information in Table 6 to divide the entropy production rate $\partial_{\mu}s^{\mu}$ into the following six terms,
$\partial_{\mu}s^{\mu}=\partial_{\mu}s^{\mu}_{S.S}+\partial_{\mu}s^{\mu}_{PS.PS}+\partial_{\mu}s^{\mu}_{V.V}+\partial_{\mu}s^{\mu}_{PV.PV}+\partial_{\mu}s^{\mu}_{T.T}+\partial_{\mu}s^{\mu}_{PT.PT}\;.$ (27)
Here, $\partial_{\mu}s^{\mu}_{S.S}$ and $\partial_{\mu}s^{\mu}_{PS.PS}$ denote the contributions bilinear in scalar and pseudo-scalar fields, respectively, and other terms are labeled in a similar way, with the correspondences $S$: scalar, $PS$: pseudo-scalar, $V$: vector, $PV$: pseudo-vector, $T$: tensor, and $PT$: pseudo-tensor. Explicitly, one can write each of these contributions in matrix form as
$\partial_{\mu}s_{S.S}^{\mu}$ $=\frac{1}{T^{2}}\begin{pmatrix}\theta\\ TD\beta\\ D\zeta\\ \mathcal{Z}\end{pmatrix}^{\text{T}}\left(\begin{array}[]{cccc}h_{2}&h_{4}/2&k_{2}/2&(2m_{3}-t_{2})/4\\ h_{4}/2&0&0&-t_{1}/4\\ k_{2}/2&0&0&-u_{1}/4\\ (2m_{3}-t_{2})/4&-t_{1}/4&-u_{1}/4&-w_{1}/2\\ \end{array}\right)\begin{pmatrix}\theta\\ TD\beta\\ D\zeta\\ \mathcal{Z}\end{pmatrix}\;,$ (32)
$\partial_{\mu}s_{PS.PS}^{\mu}$ $=-\frac{1}{2T^{2}}w_{2}\mathcal{H}^{2}\;,$ (33)
$\partial_{\mu}s_{V.V}^{\mu}$ $=-\frac{1}{T^{2}}\begin{pmatrix}\mathcal{J}_{h}^{\mu}\\ \nabla^{\mu}\zeta\\ \mathcal{E}^{\mu}\\ \mathcal{G}^{\mu}\\ \mathcal{X}^{\mu}\end{pmatrix}^{\text{T}}\left(\begin{array}[]{ccccc}-h_{1}&(g_{1}-k_{1})/2&(l_{1}-n_{1})/2&(2m_{5}-t_{3})/4&m_{1}/2\\ (g_{1}-k_{1})/2&i_{1}&(p_{1}-o_{1})/2&-(2f_{3}+u_{2})/4&-f_{1}/2\\ (l_{1}-n_{1})/2&(p_{1}-o_{1})/2&q_{1}&(2r_{3}-v_{3})/4&r_{1}/2\\ (2m_{5}-t_{3})/4&-(2f_{3}+u_{2})/4&(2r_{3}-v_{3})/4&-w_{9}/2&-w_{7}/4\\ m_{1}/2&-f_{1}/2&r_{1}/2&-w_{7}/4&0\\ \end{array}\right)\begin{pmatrix}\mathcal{J}^{h}_{\mu}\\ \nabla_{\mu}\zeta\\ \mathcal{E}_{\mu}\\ \mathcal{G}_{\mu}\\ \mathcal{X}_{\mu}\end{pmatrix}\;,$ (39)
$\partial_{\mu}s_{PV.PV}^{\mu}$ $=-\frac{1}{T^{2}}\begin{pmatrix}\mathcal{B}^{\mu}\\ \mathcal{T}^{\mu}\\ \mathcal{Y}^{\mu}\end{pmatrix}^{\text{T}}\left(\begin{array}[]{ccc}q_{2}&-(2r_{8}+v_{1})/4&r_{6}/2\\ -(2r_{8}+v_{1})/4&-w_{19}/2&-w_{6}/4\\ r_{6}/2&-w_{6}/4&0\\ \end{array}\right)\begin{pmatrix}\mathcal{B}^{\mu}\\ \mathcal{T}^{\mu}\\ \mathcal{Y}^{\mu}\end{pmatrix}\;,$ (43)
$\partial_{\mu}s_{T.T}^{\mu}$ $=\frac{1}{T^{2}}\begin{pmatrix}\sigma^{\mu\nu}\\ \mathcal{F}^{\mu\nu}_{S}\end{pmatrix}^{\text{T}}\left(\begin{array}[]{ccc}h_{3}&(2m_{10}-t_{7})/4\\ (2m_{10}-t_{7})/4&-w_{10}/2\\ \end{array}\right)\begin{pmatrix}\sigma^{\mu\nu}\\ \mathcal{F}^{\mu\nu}_{S}\end{pmatrix}\;,$ (46)
$\partial_{\mu}s_{PT.PT}^{\mu}$ $=\frac{1}{4T^{2}}w_{24}\mathcal{I}^{\mu\nu}_{S}\mathcal{I}_{\mu\nu\,S}\;.$ (47)
The semi-positivity of $\partial_{\mu}s^{\mu}$ requires each of these parts to be semi-positive. There are two equivalent approaches to establish semi-positivity: (1) diagonalize the relevant matrices such that the results are expressed in terms of perfect square formulas, or (2) use the fact that all eigenvalues of a matrix are non-negative if and only if all principal minors of the matrix are non-negative. In this work we use the second approach, which leads to the following conditions,
$\partial_{\mu}s^{\mu}_{S.S}\geq 0\quad\implies\quad$$\begin{cases}&h_{2}\geq 0,\quad w_{1}\leq 0,\quad h_{4}=k_{2}=t_{1}=u_{1}=0,\\ &-8h_{2}w_{1}\geq(2m_{3}-t_{2})^{2},\end{cases}$ (48)
$\partial_{\mu}s^{\mu}_{PS.PS}\geq 0\quad\implies\quad$ $\hskip 19.91684ptw_{2}\leq 0,$ (49)
$\partial_{\mu}s^{\mu}_{V.V}\geq 0\quad\implies\quad$$\begin{cases}&{m_{1}=f_{1}=r_{1}=w_{7}=0},\quad h_{1}\leq 0,~{}i_{1}\geq 0,~{}w_{9}\leq 0,\\ &q_{1}+h_{1}=0,\quad g_{1}-o_{1}=0,\quad t_{3}+v_{3}=0,\\ &-4h_{1}i_{1}\geq(k_{1}-g_{1})^{2},\quad 8h_{1}w_{9}\geq(2m_{5}-t_{3})^{2},\quad-8i_{1}w_{9}\geq(2f_{3}+u_{2})^{2},\\ &h_{1}(2f_{3}+u_{2})^{2}-i_{1}(2m_{5}-t_{3})^{2}+2w_{9}(k_{1}-g_{1})^{2}\\ &\hskip 19.91684pt+(2f_{3}+u_{2})(2m_{5}-t_{3})(k_{1}-g_{1})+8h_{1}i_{1}w_{9}\geq 0,\end{cases}$ (50)
$\partial_{\mu}s^{\mu}_{PV.PV}\geq 0\quad\implies\quad$$\begin{cases}&{q_{2}\geq 0},\quad{w_{19}\leq 0},\quad{r_{6}=w_{6}=0},\\ &-8q_{2}w_{19}\geq(2r_{8}+v_{1})^{2},\end{cases}$ (51)
$\partial_{\mu}s^{\mu}_{T.T}\geq 0\quad\implies\quad$$\begin{cases}&{h_{3}\geq 0},\quad{w_{10}\leq 0},\\ &-8h_{3}w_{10}\geq(2m_{10}-t_{7})^{2},\end{cases}$ (52)
$\partial_{\mu}s^{\mu}_{PT.PT}\geq 0\quad\implies\quad$ $\hskip 19.91684pt{w_{24}\geq 0}.$ (53)
These constraints include 13 equalities, which reduce the number of independent coefficients to $36-13=23$. The remaining quantities are further constrained by the inequalities introduced in Eqs. (48)–(53), which can be classified as follows: 11 inequalities demand that a certain coefficient has a definite sign, 6 inequalities set a bound for the product of two coefficients, and 1 inequality constrains the product of three coefficients. Taking all of these into account, we can construct the final form of the dissipative currents, as will be done in the following.
IV Final expressions and comparison with previous results
In this section we provide the explicit expressions for the dissipative currents at first order in gradients. The symmetric dissipative part of the energy-momentum tensor takes the following form:
$\delta T^{\mu\nu}_{S}=$ $2u^{(\mu}\left[-\kappa_{h}(\mathcal{J}_{h}^{\nu)}-\mathcal{E}^{\nu)})+\kappa_{\zeta}\,\nabla^{\nu)}\zeta+\kappa_{\mathcal{G}}\,\mathcal{G}^{\nu)}\right]+\left(\zeta_{b}\theta+\zeta_{\mathcal{Z}}\,\mathcal{Z}\right)\Delta^{\mu\nu}+2\left(\eta\sigma^{\mu\nu}+\eta_{\mathcal{F}}\,\mathcal{F}_{S}^{\mu\nu}\right),$ (54)
where the thermo-hydrodynamic fields, the corresponding transport coefficients, their relations with factors introduced in Tables 1-5, and the constraints and physical interpretations of these coefficients are given in Table 7.
Thermo-hydrodynamic field Coefficient Constraint Interpretation
$\theta=\nabla^{\mu}u_{\mu}$ $\zeta_{b}={h_{2}}/{T}$ $\zeta_{b}\geq 0$ Bulk viscosity
$\sigma^{\mu\nu}=\nabla^{(\mu}u^{\nu)}-\frac{1}{3}\theta\Delta^{\mu\nu}$ $\eta={h_{3}}/{(2T)}$ $\eta\geq 0$ Shear viscosity
$\mathcal{Z}=u^{[\mu}\nabla^{\nu]}\Omega_{\mu\nu}$ $\zeta_{\mathcal{Z}}=m_{3}/T$ $4\zeta\chi_{\mathcal{Z}}\geq(\zeta_{\mathcal{Z}}+\chi_{\theta})^{2}$ Gyro-bulk viscosity
$\mathcal{F}^{\mu\nu}_{S}={-u_{\rho}\nabla^{(\mu}\Omega^{\nu)\rho}-\frac{1}{3}\Delta^{\mu\nu}\mathcal{Z}}$ $\eta_{\mathcal{F}}={m_{10}}/{(2T)}$ $16\eta\chi_{\mathcal{F}}\geq(4\eta_{\mathcal{F}}+\chi_{\sigma})^{2}$ Gyro-shear viscosity
Table 7: Summary of contributions to $\delta T^{\mu\nu}_{S}$ in Eq. (54) from various thermo-hydrodynamic fields, along with the corresponding transport coefficients. The first three coefficients in Eq. (54) are not independent, as they also appear in $\delta T^{\mu\nu}_{A}$ in Eq. (55) due to the matching condition (13); they are listed in Table 8.
In Table 7, $\eta$ and $\zeta_{b}$ are the shear and bulk viscosities. Two new transport coefficients, $\zeta_{\mathcal{Z}}$ and $\eta_{\mathcal{F}}$, are identified for the symmetric part of the stress-energy tensor, corresponding to the gradients of the spin potential. The first three coefficients in Eq. (54) multiplying vector terms also appear in $\delta T^{\mu\nu}_{A}$, according to Eq. (55), because they are involved in the matching condition $u_{\mu}(T^{\mu\nu}_{S}+T^{\mu\nu}_{A})=0$ in Eq. (13); they have been omitted in Table 7 and quoted in Table 8.
The anti-symmetric contribution to the energy-momentum tensor reads:
$\delta T^{\mu\nu}_{A}=$ $-2u^{[\mu}\left[-\kappa_{h}(\mathcal{J}_{h}^{\nu]}-\mathcal{E}^{\nu]})+\kappa_{\zeta}\,\nabla^{\nu]}\zeta+\kappa_{\mathcal{G}}\,\mathcal{G}^{\nu]}\right]-2\gamma_{\phi}\phi^{\mu\nu}-2\gamma_{\Xi}\Xi^{\mu\nu}\;,$ (55)
and the relevant transport coefficients are summarized in Table 8. The coefficients $\kappa_{h}$ and $\gamma_{\phi}$ have been previously obtained in Ref. [7] and named boost heat conductivity and rotational viscosity, denoted as $\lambda$ and $\gamma$ respectively. It should be noted that the current associated with the coefficient $\kappa_{h}$ does not precisely match that of $\lambda$ in Ref. [7] due to the choice of the matching condition. Nevertheless, they basically have the same physical meaning. Studies on the rotational viscosity coefficient have been also conducted in Refs. [41, 42]. We thus identify three new transport coefficients: $\kappa_{\zeta}$, $\kappa_{\mathcal{G}},$ and $\gamma_{\Xi}$, which appear due to the inclusion of the gradient of the spin potential and the finite chemical potential.
Thermo-hydrodynamic field Coefficient Constraint Interpretation
$\mathcal{J}_{h}^{\mu}-\mathcal{E}^{\mu}=Du^{\mu}-\frac{1}{T}\nabla^{\mu}T+T(\Omega^{\mu\nu}-\varpi^{\mu\nu})u_{\nu}$ $\kappa_{h}=-h_{1}/(2T)$ $\kappa_{h}\geq 0$ Boost heat conductivity
$\nabla^{\mu}\zeta$ $\kappa_{\zeta}={k_{1}}/{(2T)}$ $8T\kappa_{h}\mathcal{K}_{\zeta}\geq(T\mathcal{K}_{h}-2\kappa_{\zeta})^{2}$ Chemical-boost heat conductivity
$\mathcal{G}^{\mu}=2\Delta^{\mu\tau}\nabla^{\rho}\Omega_{\rho\tau}$ $\kappa_{\mathcal{G}}=-m_{5}/(2T)$ $16\kappa_{h}\chi_{\mathcal{G}}\geq(8\kappa_{\mathcal{G}}-\chi_{h})^{2}$ Gyro-boost heat conductivity
$\phi^{\mu\nu}=T\Delta^{\mu[\sigma}\Delta^{\rho]\nu}(\Omega_{\rho\sigma}-\varpi_{\rho\sigma})$ $\gamma_{\phi}=q_{2}/T$ $\gamma_{\phi}\geq 0$ Rotational viscosity
$\Xi^{\mu\nu}=-u_{\rho}\nabla^{[\mu}\Omega^{\nu]\rho}$ $\gamma_{\Xi}={r_{8}}/{T}$ $16\gamma_{\phi}\chi_{\Xi}\geq(4\gamma_{\Xi}+\chi_{\phi})^{2}$ Gyro-rotation viscosity
Table 8: Summary of contributions to $\delta T^{\mu\nu}_{A}$ in Eq. (55) from various thermo-hydrodynamic fields, along with the corresponding transport coefficients.
For the conserved vector current, we obtain the following expression:
$\delta j^{\mu}=\mathcal{K}_{\zeta}\nabla^{\mu}\zeta+\mathcal{K}_{h}\left(\mathcal{J}_{h}^{\mu}-\mathcal{E}^{\mu}\right)+\mathcal{K}_{\mathcal{G}}\,\mathcal{G}^{\mu}.$ (56)
A summary of the corresponding transport coefficients, along with their respective constraints and physical interpretations, is provided in Table 9. The coefficient $\mathcal{K}_{\zeta}$ relates the dissipative current to the gradient of the chemical potential and can then be identified as the diffusion coefficient. However, in the context of relativistic spin hydrodynamics, we observe the appearance of two additional transport coefficients, $\mathcal{K}_{h}$ and $\mathcal{K}_{\mathcal{G}}$, which are not present in the traditional formulation without the spin tensor.
Thermo-hydrodynamic field Coefficient Constraint Interpretation
$\nabla^{\mu}\zeta$ $\mathcal{K}_{\zeta}={i_{1}}/{T^{2}}$ $\mathcal{K}_{\zeta}\geq 0$ Diffusion coefficient
$\mathcal{J}_{h}^{\mu}-\mathcal{E}^{\mu}=Du^{\mu}-\frac{1}{T}\nabla^{\mu}T+T(\Omega^{\mu\nu}-\varpi^{\mu\nu})u_{\nu}$ $\mathcal{K}_{h}={g_{1}}/{T^{2}}$ $8T\kappa_{h}\mathcal{K}_{\zeta}\geq(T\mathcal{K}_{h}-2\kappa_{\zeta})^{2}$ Diffusion-boost heat conductivity
$\mathcal{G}^{\mu}=2\Delta^{\mu\tau}\nabla^{\rho}\Omega_{\rho\tau}$ $\mathcal{K}_{\mathcal{G}}=-{f_{3}}/{T^{2}}$ $8T\chi_{G}\mathcal{K}_{\zeta}\geq(2T\mathcal{K}_{\mathcal{G}}+\chi_{\zeta})^{2}$ Diffusion-gyro-boost heat conductivity
Table 9: Summary of various transport coefficients for $\delta j^{\mu}$ in Eq. (56).
Finally, we report the expression for the spin dissipative current:
$\delta S^{\mu\lambda\nu}=$ $\frac{1}{T}\left[\Delta^{\mu[\lambda}\left(\chi_{h}(\mathcal{J}_{h}^{\nu]}-\mathcal{E}^{\nu]})+\chi_{\zeta}\nabla^{\nu]}\zeta+\chi_{\mathcal{G}}\mathcal{G}^{\nu]}\right)+\Delta^{\mu[\lambda}u^{\nu]}\left(\chi_{\theta}\theta+\chi_{\mathcal{Z}}\mathcal{Z}\right)+\left(\chi_{\sigma}\sigma^{\mu[\lambda}u^{\nu]}+\chi_{\mathcal{F}}\mathcal{F}_{S}^{\mu[\lambda}u^{\nu]}\right)\right.$
$\left.+\left(\chi_{\phi}\phi^{\mu[\lambda}u^{\nu]}+\chi_{\Xi}\Xi^{\mu[\lambda}u^{\nu]}\right)+\left(\chi_{\mathcal{H}}\epsilon^{\mu\lambda\nu\rho}u_{\rho}\mathcal{H}+\chi_{\mathcal{I}}\mathcal{I}^{\mu}_{\rho,S}\epsilon^{\lambda\nu\rho\gamma}u_{\gamma}\right)\right]\;,$ (57)
with transport coefficients listed in Table 10. The coefficients $\chi_{Z},\chi_{F},\chi_{\Xi}$ were identified in Refs. [18, 29, 35]; most coefficients related to the derivatives of the spin potential were obtained in Refs. [8, 36]. In addition to the constraints in Tables 7-10, we have another requirement:
$16T\kappa_{h}\mathcal{K}_{\zeta}\chi_{\mathcal{G}}\geq T\mathcal{K}_{\zeta}(8\kappa_{\mathcal{G}}-\chi_{h})^{2}+2\kappa_{h}(2T\mathcal{K}_{\mathcal{G}}+\chi_{\zeta})^{2}+2\chi_{\mathcal{G}}(2\chi_{\zeta}-T\mathcal{K}_{h})^{2}-(8\kappa_{\mathcal{G}}-\chi_{h})(2T\mathcal{K}_{\mathcal{G}}+\chi_{\zeta})(2\chi_{\zeta}-T\mathcal{K}_{h})\;.$ (58)
Thermo-hydrodynamic field Coefficient Constraint Interpretation
$(\mathcal{J}_{h}^{\nu}-\mathcal{E}^{\nu})=Du^{\nu}-\frac{1}{T}\nabla^{\nu}T+T(\Omega^{\nu\lambda}-\varpi^{\nu\lambda})u_{\lambda}$ $\chi_{h}=-t_{3}/T$ $16\kappa_{h}\chi_{\mathcal{G}}\geq(8\kappa_{\mathcal{G}}-\chi_{h})^{2}$ Spin-boost heat conductivity
$\nabla^{\nu}\zeta$ $\chi_{\zeta}=-u_{2}/T$ $8T\chi_{\mathcal{G}}\mathcal{K}_{\zeta}\geq(2T\mathcal{K}_{\mathcal{G}}+\chi_{\zeta})^{2}$ Spin-chemical-boost heat conductivity
$\mathcal{G}^{\nu}=2\Delta^{\nu\tau}\nabla^{\rho}\Omega_{\rho\tau}$ $\chi_{\mathcal{G}}=-w_{9}/T$ $\chi_{\mathcal{G}}\geq 0$ Spin-gyro-boost heat conductivity
$\theta=\nabla^{\mu}u_{\mu}$ $\chi_{\theta}=-t_{2}/(2T)$ $4\zeta\chi_{\mathcal{Z}}\geq(\zeta_{\mathcal{Z}}+\chi_{\theta})^{2}$ Spin-bulk viscosity
$\mathcal{Z}=u^{[\mu}\nabla^{\nu]}\Omega_{\mu\nu}$ $\chi_{\mathcal{Z}}=-w_{1}/(2T)$ $\chi_{\mathcal{Z}}\geq 0$ Spin-gyro-bulk viscosity
$\sigma^{\mu\lambda}=\nabla^{(\mu}u^{\lambda)}-\frac{1}{3}\theta\Delta^{\mu\lambda}$ $\chi_{\sigma}=-t_{7}/T$ $16\eta\chi_{\mathcal{F}}\geq(4\eta_{\mathcal{F}}+\chi_{\sigma})^{2}$ Spin-shear viscosity
$\mathcal{F}^{\mu\lambda}_{S}=-u_{\rho}\nabla^{(\mu}\Omega^{\lambda)\rho}-\frac{1}{3}\Delta^{\mu\lambda}\mathcal{Z}$ $\chi_{\mathcal{F}}=-w_{10}/T$ $\chi_{\mathcal{F}}\geq 0$ Spin-gyro-shear viscosity
$\phi^{\mu\lambda}=T\Delta^{\mu[\sigma}\Delta^{\rho]\lambda}(\Omega_{\rho\sigma}-\varpi_{\rho\sigma})$ $\chi_{\phi}=2v_{1}/T$ $16\gamma_{\phi}\chi_{\Xi}\geq(4\gamma_{\Xi}+\chi_{\phi})^{2}$ Spin-rotation viscosity
$\Xi^{\mu\lambda}=-u_{\rho}\nabla^{[\mu}\Omega^{\lambda]\rho}$ $\chi_{\Xi}=-2w_{19}/T$ $\chi_{\Xi}\geq 0$ Spin-gyro-rotation viscosity
$\mathcal{H}=\epsilon^{\rho\sigma\lambda\gamma}u_{\lambda}\partial_{\gamma}\Omega_{\rho\sigma}$ $\chi_{\mathcal{H}}=-w_{2}/T$ $\chi_{\mathcal{H}}\geq 0$ Spin-gyro-pseudo bulk viscosity
$\mathcal{I}^{\mu\rho}_{S}=u_{\gamma}\epsilon^{\theta\lambda\gamma(\mu}\nabla^{\rho)}\Omega_{\theta\lambda}-\Delta^{\mu}_{\rho}\mathcal{H}/3$ $\chi_{\mathcal{I}}=w_{24}/(2T)$ $\chi_{\mathcal{I}}\geq 0$ Spin-gyro-pseudo shear viscosity
Table 10: Summary of various transport coefficients for $\delta S^{\mu\lambda\nu}$ in Eq. (57).
It is worth mentioning that most of the references cited in this section follow a similar approach in deriving the form of dissipative currents, assuming the validity of the following local thermodynamic relations in spin hydrodynamics:
$\begin{split}&Ts+\mu n=\rho+p-\frac{1}{2}\omega_{\mu\nu}S^{\mu\nu}\;,\\ &{\rm d}p=s\,{\rm d}T+n\,{\rm d}\mu+\frac{1}{2}S^{\mu\nu}{\rm d}\omega_{\mu\nu}\;,\end{split}$ (59)
where $\mu\equiv\zeta T$ is the chemical potential, $\rho$ the proper energy density, $p$ the pressure, $n$ the charge density, $S^{\mu\nu}\equiv u_{\lambda}{\cal S}^{\lambda,\mu\nu}$ is the “spin density” and $\omega_{\mu\nu}=T\Omega_{\mu\nu}$. Nevertheless, the differential relation in eq. (59) may not be fulfilled, as it was pointed out in Ref. [39], so that using the equation (4) as a starting point appears to be a safer and more general method.
V Application to quantum kinetic theory
Since the second law of thermodynamics is a macroscopic statement, any (first-order) microscopic theory has to comply with it. One such example is the formulation of spin hydrodynamics from quantum kinetic theory [26, 43, 44, 45]. Before being able to properly apply the results derived in the preceding sections, we have to clarify the power-counting used in quantum kinetic theory, which is referred to as the $\hbar$-gradient expansion and denotes an expansion around the classical limit. In particular, in the standard counting that treats spin-related quantities as small quantum corrections, one has $\hbar\partial_{\lambda}S^{\lambda\mu\nu}\sim T_{A}^{\mu\nu}\sim\mathcal{O}(\hbar^{2})$. On the other hand, $j^{\mu}\sim T_{S}^{\mu\nu}\sim\mathcal{O}(1)$ are of zeroth order, with corrections emerging at second order. These corrections have not been computed in Refs. [26, 43, 44, 45], so we are a bit limited in our analysis. We also have to revise the conditions given in Eqs. (14) and (15). The reason is that, since in the aforementioned references the currents $\delta j^{\mu}$ and $\delta T_{S}^{\mu\nu}$ are only evaluated up to first order in $\hbar$, the matching conditions are $u_{\mu}\delta j^{(0)\mu}=0$ and $u_{\mu}\delta T_{S}^{(0)\mu\nu}=0$. Here, quantities marked with an index $(0)$ denote the contributions of zeroth order in quantum corrections. The form of the currents is then
$\delta j^{\mu}$ $=\mathcal{K}_{\zeta}^{(0)}\nabla^{\mu}\zeta+\delta j^{(2)\mu}\;,$ (60a)
$\delta T_{S}^{\mu\nu}$ $=\Delta^{\mu\nu}\zeta_{b}^{(0)}\theta+2\eta^{(0)}\sigma^{\mu\nu}+\delta T_{S}^{(2)\mu\nu}\;,$ (60b)
$\delta T_{A}^{\mu\nu}$ $=2u^{[\mu}\left(\kappa_{h}\mathcal{J}_{h}^{\nu]}-f_{2}\mathcal{E}^{\nu]}+f_{4}\mathcal{X}^{\nu]}-\kappa_{\zeta}\nabla^{\nu]}\zeta-\kappa_{\mathcal{G}}\mathcal{G}^{\nu]}\right)+\epsilon^{\mu\nu\alpha\beta}u_{\alpha}\left(-\gamma_{\phi}\mathcal{B}_{\beta}+\gamma_{\Xi}\mathcal{T}_{\beta}\right)\;,$ (60c)
$\delta S^{\lambda\mu\nu}$ $=\frac{1}{T}\Big{[}-\frac{1}{2}u^{[\mu}\epsilon^{\nu]\lambda\alpha\beta}u_{\alpha}\left(-\chi_{\phi}\mathcal{B}_{\beta}+\chi_{\Xi}\mathcal{T}_{\beta}\right)+\Delta^{\lambda[\mu}u^{\nu]}\left(k_{1}TD\beta+\chi_{\theta}\theta+\chi_{\mathcal{Z}}\mathcal{Z}+k_{4}D\zeta\right)+\chi_{\sigma}\sigma^{\lambda[\mu}u^{\nu]}+\chi_{\mathcal{F}}\mathcal{F}_{S}^{\lambda[\mu}u^{\nu]}$
$\quad+\Delta^{\lambda[\mu}\left(\chi_{h}\mathcal{J}_{h}^{\nu]}-m_{2}\mathcal{E}^{\nu]}+m_{3}\mathcal{X}^{\nu]}+\chi_{\zeta}\nabla^{\nu]}\zeta+\chi_{\mathcal{G}}\mathcal{G}^{\nu]}\right)+\chi_{\mathcal{H}}\epsilon^{\lambda\mu\nu\alpha}u_{\alpha}\mathcal{H}+\chi_{\mathcal{I}}\mathcal{I}_{S,\alpha}{}^{\lambda}\epsilon^{\mu\nu\alpha\rho}u_{\rho}\Big{]}\;.$ (60d)
The quantities $\delta j^{(2)\mu}$ and $\delta T_{S}^{(2)\mu\nu}$ are not further constrained. Compared to Eqs. (54)–(57), there are a few differences, which are due to the different matching. In particular, in Eq. (60c), there are two extra coefficients, which in the matching discussed in the previous sections become $f_{1}=\kappa_{h}$ and $f_{2}=0$, respectively. Similarly, in Eq. (60d), one would have $k_{1}=k_{2}=0$, $m_{2}=\chi_{h}$, and $m_{3}=0$. The full set of inequalities that the coefficients have to fulfill in this case in order for the entropy production to be positive at any order in quantum corrections becomes hardly tractable. However, we can read off a set of necessary (though not sufficient) conditions,
$\mathcal{K}_{\zeta}^{(0)}$ $\geq 0\;,\quad\zeta^{(0)}_{b}\geq 0\;,\quad\eta^{(0)}\geq 0\;,\quad f_{2}\geq 0\;,\quad\gamma_{\phi}\geq 0\;,\quad\chi_{\Xi}\geq 0\;,$
$\chi_{\mathcal{Z}}$ $\geq 0\;,\quad\chi_{\mathcal{F}}\geq 0\;,\quad\chi_{\mathcal{G}}\geq 0\;,\quad\chi_{\mathcal{H}}\geq 0\;,\quad\chi_{\mathcal{I}}\geq 0\;,$ (61)
that is identical to a subset of entries in Tables 710, namely those that require certain coefficients to be nonnegative. Even though we do not have access to $\delta j^{(2)\mu}$ and $\delta T_{S}^{(2)\mu\nu}$, we can still check the inequalities (V) explicitly.
For the comparison, we focus on the formulation derived in Ref. [45]. Therein, the information on dissipation is included in the bulk viscous pressure $\Pi$, the diffusion current $n^{\mu}$, the shear-stress tensor $\pi^{\mu\nu}$, and a traceless rank-two tensor $\mathfrak{t}^{\mu\nu}$ appearing in the spin tensor. These quantities are dynamical and follow relaxation-type equations [detailed in Eq. (2) of Ref. [45]]. To find their Navier-Stokes values, i.e., relate them algebraically to the first- order quantities at our disposal, we only keep those terms in the respective equations of motion that are linear in the first-order tensors. [^2] In the context of this work, we also treat the components of $\partial^{\lambda}\Omega^{\mu\nu}$ as being of first order when finding the Navier-Stokes limit of $\mathfrak{t}^{\mu\nu}$. Also, we identified $\mathcal{F}_{S}^{\mu\nu}\simeq\nabla^{\langle\mu}\kappa^{\nu\rangle}$. In particular, keeping to the notation of Ref. [45] and denoting potentially ambiguous quantities with a subscript “kin”, we have
$\Pi=-\zeta_{\text{kin}}\theta\;,\quad n^{\mu}=\kappa_{\text{kin}}\nabla^{\mu}\zeta\;,\quad\pi^{\mu\nu}=2\eta_{\text{kin}}\sigma^{\mu\nu}\;,\quad\mathfrak{t}^{\mu\nu}=\frac{1}{T}\mathfrak{d}\sigma^{\mu\nu}+\ell_{\mathfrak{t}\kappa}\mathcal{F}_{S}^{\mu\nu}\;.$ (62)
Then, the form of the currents (in the Navier-Stokes limit) is given by
$\delta j^{\mu}_{\text{kin}}$ $=\kappa_{\text{kin}}\nabla^{\mu}\zeta+\delta j^{(2)\mu}_{\text{kin}}\;,$ (63a)
$\delta T^{\mu\nu}_{S,\text{kin}}$ $=\zeta_{\text{kin}}\theta\Delta^{\mu\nu}+2\eta_{\text{kin}}\sigma^{\mu\nu}+\delta T_{S,\text{kin}}^{(2)\mu\nu}\;,$ (63b)
$\delta T^{\mu\nu}_{A,\text{kin}}$ $=2u^{[\mu}\left(\frac{\Gamma^{(a)}}{2}\frac{1}{T}\mathcal{J}_{h}^{\nu]}-\frac{1}{T}\Gamma^{(\kappa)}\mathcal{E}^{\nu]}+\Gamma^{(I)}\nabla^{\nu]}\zeta\right)-\frac{1}{2T}\Gamma^{(\omega)}\epsilon^{\mu\nu\alpha\beta}u_{\alpha}\mathcal{B}_{\beta}\;,$ (63c)
$\delta S^{\lambda\mu\nu}_{\text{kin}}$ $=\frac{2\sigma}{mT}\mathfrak{d}\sigma^{\lambda[\mu}u^{\nu]}+\frac{2\sigma}{m}\ell_{\mathfrak{t}\kappa}\mathcal{F}_{S}^{\lambda[\mu}u^{\nu]}\;,$ (63d)
with $\delta j^{(2)\mu}_{\text{kin}}$ and $\delta T_{S,\text{kin}}^{(2)\mu\nu}$ denoting the second-order contributions that have not been computed in Ref. [45]. Furthermore, $\sigma$ is the spin of the particle, and $I_{nq}$, $J_{nq}$ are thermodynamic integrals, which for Boltzmann statistics read
$I_{nq}=J_{nq}=\frac{ge^{\alpha}}{(2q+1)!!}\frac{T^{n+2}}{2\pi^{2}}z^{n+2}\sum_{j=0}^{q+1}(-1)^{j}\binom{q+1}{j}\mathrm{Ki}_{2j-2-n}(z)\;,$ (64)
with $z=m/T$ being the ratio of mass and temperature, $g=2\sigma+1$ the degeneracy factor, and $\mathrm{Ki}_{n}(z)$ the Bickley function.
Comparing Eqs. (60) and (63), we can identify
$\mathcal{K}^{(0)}_{\zeta}$ $\equiv\kappa_{\text{kin}}\;,\;\;\zeta^{(0)}_{b}\equiv\zeta_{\text{kin}}\;,\;\;\eta^{(0)}\equiv\eta_{\text{kin}}\;,$
$\kappa_{h}$ $\equiv\frac{\Gamma^{(a)}}{2T}\;,\;\;f_{2}\equiv\frac{\Gamma^{(\kappa)}}{T}\;,\;\;\kappa_{\zeta}\equiv-\Gamma^{(I)}\;,\;\;\gamma_{\phi}\equiv\frac{\Gamma^{(\omega)}}{2T}\;,\;\;\chi_{\sigma}\equiv\frac{2\sigma}{m}\mathfrak{d}\;,\;\;\chi_{\mathcal{F}}\equiv\frac{2\sigma T}{m}\ell_{\mathfrak{t}\kappa}\;,$
$f_{4}$ $=\kappa_{\mathcal{G}}=\gamma_{\Xi}=\chi_{\phi}=\chi_{\Xi}=k_{1}=\chi_{\theta}=\chi_{\mathcal{Z}}=k_{4}=\chi_{h}=m_{2}=m_{3}=\chi_{\zeta}=\chi_{\mathcal{G}}=\chi_{\mathcal{H}}=\chi_{\mathcal{I}}\equiv 0\;.$ (65)
We emphasize that the fact that many coefficients vanish is a consequence of the simple truncation employed in Sec. IX of Ref. [45]. When using a higher-order truncation, a lot of the coefficients will become nonzero, allowing for further nontrivial inequalities to be checked. These investigations are left for future work. Considering the subset of coefficients that are nonzero, the inequalities (V) become
$\kappa_{\text{kin}}\geq 0\;,\quad\zeta_{\text{kin}}\geq 0\;,\quad\eta_{\text{kin}}\geq 0\;,\quad\Gamma^{(\kappa)}\geq 0\;,\quad\Gamma^{(\omega)}\geq 0\;,\quad\ell_{\mathfrak{t}\kappa}\geq 0\;.$ (66)
The first three inequalities are well-known from the case of standard hydrodynamics, while the following three appear only for nonzero spin. Note that the conditions $\Gamma^{(\kappa)}\geq 0$ and $\Gamma^{(\omega)}\geq 0$ can be related to the damping of the spin waves, i.e., the rate at which the spin potential approaches the thermal vorticity [44].
While a formal proof for any interaction is outside the scope of this paper, we can check the inequalities (66) for a specific microscopic interaction, which we choose to be a four-fermi one, $\mathcal{L}_{\text{int}}=G(\overline{\psi}\psi)^{2}$. As one can see in Fig. 2, all coefficients are positive and the inequalities are indeed fulfilled.
$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$$10^{-8}$$10^{-6}$$10^{-4}$$10^{-2}$$10^{0}$$z$$\eta_{\text{kin}}[1/(TG^{2})]$$\kappa_{\text{kin}}[1/(T^{2}G^{2})]$$\zeta_{\text{kin}}[1/(TG^{2})]$$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$$10^{-8}$$10^{-6}$$10^{-4}$$10^{-2}$$10^{0}$$z$$z^{2}\Gamma^{(\kappa)}[G^{2}T^{8}]$$z^{2}\Gamma^{(\omega)}[G^{2}T^{8}]$$\ell_{\mathfrak{t}\kappa}[1/(TG^{2})]$
Figure 2: Left: The transport coefficients that have to be semipositive also in hydrodynamics without spin. Right: The additional coefficients that have to be semipositive in spin hydrodynamics. Here, $G$ denotes the coupling strengh of the four-fermi interaction. Note that we multiplied the coefficients $\Gamma^{(\kappa)}$ and $\Gamma^{(\omega)}$ by $z^{2}$ to render the result finite.
VI Summary and outlook
In summary, we have derived expressions for the first order dissipative corrections to the energy-momentum tensor, conserved vector current, and spin tensor incorporating the effects of spin potential gradients and finite chemical potential. The main results of this work are Eqs. (54)-(57) and the corresponding transport coefficients, which are reported, named, and described in Tables 7-10. In our derivation, we have expanded the tensor coefficients multiplying the gradients of the thermo-hydrodynamic fields from global thermodynamic equilibrium with vanishing thermal vorticity, thus taking advantage of rotational symmetry in the local rest frame.
With this foundation, one can formulate the evolution equations for the energy-momentum tensor, conserved vector current and spin tensor at first order in gradients, commonly referred to as the relativistic Navier-Stokes limit. However, the presence of 23 transport coefficients, which have to be either extracted from a suitable microscopic theory or fitted to data, may pose considerable challenges for future numerical studies, unless some of the gradients are neglected.
It is important to note that the names assigned to each transport coefficient in Tables 7-10 are based on the best available physical understanding of their roles. A deeper comprehension of the underlying mechanisms would further refine our interpretation of each new coefficient.
Acknowledgements
A.D. thanks the Department of Physics, University of Florence and INFN for the hospitality. A.D acknowleges technical discussions with Tomoki Goda and Radoslaw Ryblewski. A.D. acknowledges the financial support provided by the Polish National Agency for Academic Exchange NAWA under the Programme STER–Internationalisation of doctoral schools, Project no. PPI/STE/2020/1/00020 and the Polish National Science Centre Grants No.2018/30/E/ST2/00432. D.W. acknowledges support by the project PRIN2022 Advanced Probes of the Quark Gluon Plasma funded by ”Ministero dell’Università e della Ricerca”, the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the CRC-TR 211 “Strong-interaction matter under extreme conditions” – project number 315477589 – TRR 211, and by the State of Hesse within the Research Cluster ELEMENTS (Project ID 500/10.006).
Appendix A Decomposition of tensors into irreducible components under the rotation group
In this section, we outline the method used to decompose the tensor coefficients $H^{\mu\nu\rho\sigma},K^{\mu\nu\rho},\dots,W^{\mu\lambda\nu\rho\sigma\tau}$ in equation (11) in terms of the rotationally invariant components reported in Sec. (II). This method employs the irreducible representations of the rotation group $\mathrm{SO}(3,\mathbb{R})$.
A four-vector $V^{\mu}$ can be decomposed into two irreducible components under the rotation group associated to the space orthogonal to $u$: a scalar, obtained by contracting $V$ with $u$ and a vector with vanishing component along $u$. We will denote this decomposition of $V$ onto the irreducible representations $j=0$ and $j=1$ of $\mathrm{SO}(3,\mathbb{R})$ that is scalar and the vector representations - by writing:
$V^{\mu}:(0\oplus 1)\to(u^{\mu}\oplus\Delta^{\mu}_{\alpha})$ (67)
Note that we associate the four-vector $u^{\mu}$ to the scalar representation and the projector $\Delta^{\mu}_{\alpha}$ to the vector representation, where the extra index $\alpha$ signifies that the vector $V^{\mu}$ is obtained by applying the projector $\Delta^{\mu}_{\alpha}$ to another vector. Similarly, symmetric tensors $S^{\mu\nu}$ and antisymmetric tensors $A^{\mu\nu}$ can be expressed as follows:
$S^{\mu\nu}:(0\oplus 0\oplus 1\oplus 2)\to(u^{\mu}u^{\nu}\oplus\Delta^{\mu\nu}\oplus u^{(\mu}\Delta^{\nu)}_{\alpha}\oplus\Delta^{(\mu}_{\alpha}\Delta^{\nu)}_{\beta}-\frac{1}{3}\Delta_{\alpha\beta}\Delta^{\mu\nu}),$ (68)
$A^{\mu\nu}:(1\oplus 1)\to(u^{[\mu}\Delta^{\nu]\alpha}\oplus\epsilon^{\mu\nu\tau\alpha}u_{\tau}).$
where the extra indices $\alpha$ and $\beta$ again signify that tose expressions are projectors (or proportional to projectors) to be applied to tensors. The $2$ in eq. (68) denotes the irreducible 5-dimensional representation pertaining to Euclidean symmetric traceless tensors. Note that the anti-symmetric tensor requires the use of the Levi-Civita tensor to generate an independent vector projector.
By using the above decompositions as building blocks, we can decompose tensors of any rank into irreducible components under the rotation group. For example, let us consider the gradient of the spin potential $\partial_{\rho}\Omega_{\lambda\nu}$, which is a rank-3 tensor antisymmetric in the last 2 indices. We can then write, by using the decompositions in (67) and (68):
$\partial_{\rho}\Omega_{\sigma\tau}:(0\oplus 1)\otimes(1\oplus 1)\to(u_{\rho}\oplus\Delta_{\rho\alpha})\otimes\left(\frac{1}{2}(u_{\sigma}\Delta_{\tau\beta}-u_{\tau}\Delta_{\sigma\beta})\oplus\epsilon_{\sigma\tau\gamma\beta}u^{\gamma}\right).$ (69)
where the first factor $(0\oplus 1)$ refers to the gradient vector $\partial_{\rho}$ and the second, i.e. $(1\oplus 1)$, to the spin potential tensor $\Omega$. By expanding the above tensor product and taking into account the known rules for the decomposition into irreducible representations, we get:
$(0\otimes 1)=1\to u_{\rho}u_{[\sigma}\Delta_{\tau]\beta}$
$(0\otimes 1)=1\to u_{\rho}\epsilon_{\sigma\tau\gamma\beta}u^{\gamma}$
$(1\otimes 1)\to\Delta_{\rho}^{\alpha}u_{[\sigma}\Delta_{\tau]\beta}$
$(1\otimes 1)\to\Delta^{\alpha}_{\rho}\epsilon_{\sigma\tau\gamma\beta}u^{\gamma}$ (70)
We now can reduce the ($1\otimes 1$) representation into $j=0,1,2$ and associate to each a corresponding projector. In order to get the correct one for the scalar representation $j=0$, one has to contract the indices $\alpha,\beta$ in the (A); for the vector representation $j=1$ anti-symmetrize them by using a Levi-Civita projector $\epsilon^{\alpha\beta\lambda\,\delta}u_{\lambda}$ and for the tensor representation $j=2$ to symmetrize the indices $\alpha,\beta$ and remove the trace. Therefore:
$(1\otimes 1)\to\Delta_{\rho}^{\alpha}u_{[\sigma}\Delta_{\tau]\beta}=\begin{cases}0\to u_{[\sigma}\Delta_{\tau]\rho}\\ 1\to\epsilon^{\alpha\beta\lambda\,\delta}u_{\lambda}\Delta_{\rho\alpha}u_{[\sigma}\Delta_{\tau]\beta}\\ 2\to\Delta_{\rho(\alpha}u_{[\sigma}\Delta_{\tau]\beta)}-\frac{1}{3}\Delta_{\alpha\beta}u_{[\sigma}\Delta_{\tau]\rho}\end{cases}$
$(1\otimes 1)\to\Delta^{\alpha}_{\rho}\epsilon_{\sigma\tau\gamma\beta}u^{\gamma}=\begin{cases}0\to\epsilon_{\sigma\tau\gamma\rho}u^{\gamma}\\ 1\to\epsilon^{\alpha\beta\lambda\,\delta}u_{\lambda}\Delta_{\rho\alpha}\epsilon_{\sigma\tau\gamma\beta}u^{\gamma}\\ 2\to\Delta_{\rho(\alpha}\epsilon_{\sigma\tau\gamma\beta)}u^{\gamma}-\frac{1}{3}\Delta_{\alpha\beta}\epsilon_{\sigma\tau\gamma\rho}u^{\rho}\end{cases}$
Altogether, $\partial_{\rho}\Omega_{\lambda\nu}$ can be decomposed into 8 irreducible components under rotation and written as:
$\partial_{\rho}\Omega_{\sigma\tau}=$ $c_{1}\mathcal{X}^{\gamma}u_{\rho}u_{[\sigma}\Delta_{\tau]\gamma}+c_{2}\mathcal{Y}^{\gamma}u_{\rho}\epsilon_{\sigma\tau\lambda\gamma}u^{\lambda}+c_{3}\mathcal{Z}u_{[\sigma}\Delta_{\tau]\rho}+c_{4}\mathcal{T}^{\gamma}u_{[\sigma}\epsilon_{\rho\tau]\lambda\gamma}u^{\lambda}+c_{5}\mathcal{F}^{\alpha\beta}\Delta_{\rho(\alpha}u_{[\sigma}\Delta_{\tau]\beta)}$
$+c_{6}\mathcal{H}\epsilon_{\sigma\tau\gamma\rho}u^{\gamma}+c_{7}\mathcal{G}_{\gamma}\epsilon_{\rho}^{~{}~{}\beta\lambda\gamma}u_{\lambda}\epsilon_{\sigma\tau\alpha\beta}u^{\alpha}+c_{8}\mathcal{I}^{\alpha\beta}\Delta_{\rho(\alpha}\epsilon_{\sigma\tau\gamma\beta)}u^{\gamma}.$ (71)
where $c_{i},\ i=1,2,\dots,8$, are arbitrary constants whose setting amounts to a rescaling of the scalar, vector and tensor coefficients. Note that in the expression above we have merged the trace term of the $j=2$ representations into the $j=0$ representations, without loss of generality. The above expression corresponds to the eq. (III) in the main text with a suitable choice of the constants $c_{i},\ i=1,2,\dots,8$.
The same method can be applied to the decomposition of the tensor coefficients in the equation (11), with the remarkable simplification that we are only interested in the scalar components. For instance, the tensor $H^{\mu\nu\rho\sigma}$ has the following decomposition:
$H^{\mu\nu\rho\sigma}=S^{\mu\nu}\otimes S^{\rho\sigma}:(0\oplus 0\oplus 1\oplus 2)\otimes(0\oplus 0\oplus 1\oplus 2).$ (72)
Since:
[…]
according to the equation (72), the tensor $H^{\mu\nu\rho\sigma}$ can be decomposed as follows:
$H^{\mu\nu\rho\sigma}:2(0\oplus 0\oplus 1\oplus 2)\oplus 1\oplus 1\oplus 0\oplus 1\oplus 2\oplus 1\oplus 2\oplus 3\oplus 2\oplus 2\oplus 1\oplus 2\oplus 3\oplus 0\oplus 1\oplus 2\oplus 3\oplus 4$
As it can be inferred from the above expression, the tensor has six independent scalar components and the projectors can be built by combining in all possible ways the elementary building blocks $u,\Delta$ and $\epsilon$. These are shown in Table 1 in Sec. II.